OpenStudy (anonymous):

Having some problems doing a Differential equation, So it says dy/dx=(y*cos(x))/(1-y^2) with y(0)=1 I have gotten it past the point of integrating both sides but I am having trouble isolating the y to the left side of the equation, here is my integrated problem: y^2/2-ln(y)=-sin(x)+c ( I originally got ln|y|-y^2/2 for the left but found when checking my work with a basic integration calculator it seems to indicate that the Absolute value signs were not necessary, some concrete explanation on that would be appreciated) But i got it to the point where i have y^2-y=e^(-2sin(x)+2c)

OpenStudy (anonymous):

But continuing too much further is seeming to cause my y to cancel, Would really appreciate a little guidence so i am more sure of myself here

OpenStudy (unklerhaukus):

you have solved the DE plug in the initial value and solve for the consntant c

OpenStudy (anonymous):

Yes but Should i not isolate y before continuing? also wanted to make sure everything i did is correct\

OpenStudy (unklerhaukus):

many DE cannot be solve to the form y = ... y^2/2 - ln(y) = ... is fine

OpenStudy (anonymous):

Wow really? all this time I have had quite a few i could not isolate i always thought it was error on my part, in that case it would be..... 1^2/2 - ln(1)= -sin(0)+c which gives me a value of 1/2 for c?

OpenStudy (anonymous):

therefore y^2/2 -ln(y) = -sin(x) +1/2 ?

OpenStudy (unklerhaukus):

yep

OpenStudy (unklerhaukus):

what do you get if you take the derivative on both sides of this equation (with respect to x)?

OpenStudy (anonymous):

A bit confused by this forgive me for not understanding as I did not do a good job of associating vocabulary with the methodology of this chapter, i find myself doing without knowing the names of the strategies most likely i know how to do it just do not understand the term for it XD s

OpenStudy (anonymous):

Oh hold on

OpenStudy (anonymous):

i thought that said what happend if you take the implicate derivative with respect to x

OpenStudy (anonymous):

so simply isolating to X after i integrate?

OpenStudy (unklerhaukus):

it did, i changed it because the words

OpenStudy (unklerhaukus):

take this equation (your solution) y^2/2 - ln(y) = -sin(x) +1/2 take the derivative on both sides you should get something like ( f(x) ) dy/dx = g(x)

OpenStudy (anonymous):

y-1/y dy/dx=-sin(x)? been far too long since i did one of these

OpenStudy (unklerhaukus):

right, now solve for dy/dx

OpenStudy (anonymous):

Correction it should be -cos(x) on the right

OpenStudy (unklerhaukus):

oh yeah,

OpenStudy (anonymous):

so then (-ycos(x))/y-1=dy/dx

OpenStudy (anonymous):

seems liek im close but perhaps my final equation b4 deriving was not correct?

OpenStudy (unklerhaukus):

so you have (y-1/y) dy/dx = -cos(x) dy/dx = -cos(x)/(y-1/y) = -ycos(x)/(y^2-1) = -ycos(x)/(1-y^2)

OpenStudy (anonymous):

Having trouble with were that square is coming from

OpenStudy (anonymous):

I see you inversed and mult, but..

OpenStudy (unklerhaukus):

ok we had something like \[\frac{-1}{y-1/y}\] we multiply by one \[\frac{-1}{y-1/y}\times1\\=\frac{-1}{y-1/y}\times\frac yy\\=\frac{-y}{y^2-y/y}=\frac{-y}{y^2-1}\\ =\frac{y}{1-y^2}\]

OpenStudy (unklerhaukus):

dy/dx = = ycos(x)/(1-y^2) *

OpenStudy (unklerhaukus):

like the original DE

OpenStudy (anonymous):

Ahhhhh i see

OpenStudy (anonymous):

I think that about does it for that one, will most likely have a few other questions later but thank you for your help

OpenStudy (unklerhaukus):

In regard to your question about the absolute value signs in the logarithm, the different solution sets can be seen attached.

OpenStudy (anonymous):

I see thank you

OpenStudy (unklerhaukus):

y^2/2 -ln(y) = -sin(x) +1/2 y^2 - ln y^2 = -2sin(x) + 1 (maybe this is the best form for your solution)

OpenStudy (anonymous):

alright one other thing I would like to put into words why the inverse sin of 2 is not a definitive value to use to solve for in this equation, arcsin2=c equation is just an example really need to just explain arcsin2

OpenStudy (unklerhaukus):

the sine function take any real input, and returns values between -1 and 1,

OpenStudy (unklerhaukus):

the inverse sine function takes values between -1 and +1, and returns values between -π/2 and +π/2

OpenStudy (unklerhaukus):

arcsin(2) asks for the angle thats sine is 2,

OpenStudy (unklerhaukus):